# Inverse Hamiltonian reduction in type A and generalized slices in the affine Grassmannian

### v1

@@ 7.3. Equivariance of the multiplication map
(7.3.3) $\begin{split}u_{1}^{\prime}&=\begin{pmatrix}1&a_{1,1}&a_{1,2}&\dots&a_{1,k-1}&a_{1,k}\\ 0&1&a_{2,2}&\dots&a_{2,k-1}&a_{2,k}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\dots&1&a_{k,k}\\ 0&0&\dots&0&0&1\end{pmatrix}~{},\\ u_{3}^{\prime}&=\begin{pmatrix}1&c_{1,1}&c_{1,2}&\dots&c_{1,k-1}&c_{1,k}\\ 0&1&c_{2,2}&\dots&c_{2,k-1}&c_{2,k}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\dots&1&c_{k,k}\\ 0&0&\dots&0&0&1\end{pmatrix}~{},\end{split}$
@@
(7.3.6) $\begin{split}\Phi_{\alpha}(nu_{1}t^{-\alpha}u_{3}t^{\alpha})&=\Phi_{\alpha}(u_{1}t^{-\alpha}u_{3}t^{\alpha})~{},\\ \zeta_{\alpha}(nu_{1}t^{-\alpha}u_{3}t^{\alpha})&=\zeta_{\alpha}(u_{1}t^{-\alpha}u_{3}t^{\alpha})~{}.\\ \end{split}$
@@
By Lemma 7.7$(ii)$, $\mathbf{m}(y_{1},y_{2})=ny_{1}y_{2}n_{-}$ for some $n\in N^{\alpha}[t]$ and $n_{-}\in N_{-}[t]$. Therefore,
@@
But, from Proposition 7.1$(iii)$, $x_{-\alpha}(\underline{v})nx_{-\alpha}(\underline{v})^{-1}=n^{\prime}\in N^{\alpha}[t]$. Putting things together,
@@ 7.4. Inverse Hamiltonian reduction for the translation group
From Lemma 7.7$\,(iii)$, and Proposition 7.5, we see that the composition $\mathrm{Gr}^{0}_{-\alpha}\times\mathrm{Gr}_{\mu+\alpha}\xrightarrow{\bf m}\mathrm{Gr}_{\mu}\xrightarrow{\Phi_{\alpha}}\mathfrak{G}_{\alpha}$ factors through the inclusion $\overset{\circ}{\mathfrak{G}}_{\alpha}\hookrightarrow\mathfrak{G}_{\alpha}$. So, the image of the multiplication map lies inside the preimage of $\overset{\circ}{\mathfrak{G}_{\alpha}}$, i.e.,
(7.4.1) ${\bf m}:\mathrm{Gr}_{-\alpha}^{0}\times\mathrm{Gr}_{\mu+\alpha}\hookrightarrow(\overset{\circ}{\mathfrak{G}}_{\alpha})_{\mu}~{}.$
@@
This follows immediately from Lemma 7.7$(iii)$ and Proposition 7.6, since $\xi_{\alpha}$ is fully determined by the values of $\Phi_{\alpha}$ and $\zeta_{\alpha}$. ∎
So we see that $\xi_{\alpha}$ gives a putative section of ${\bf m}:\mathrm{Gr}_{-\alpha}^{0}\times\mathrm{Gr}_{\mu+\alpha}\hookrightarrow\Phi_{\alpha}^{-1}(\overset{\circ}{\mathfrak{G}_{k}^{a}})_{\mu}$ to the first factor $\mathrm{Gr}_{-\alpha}^{0}$. We want to extend this to a full section.
@@
Let $y\in(\overset{\circ}{\mathfrak{G}_{\alpha}})_{\mu}$, then
@@
(7.4.24) $\begin{split}F_{\alpha}:(\overset{\circ}{\mathfrak{G}}_{\alpha})_{\mu}&\rightarrow\mathrm{Gr}^{0}_{-\alpha}\times\mathrm{Gr}_{\mu+\alpha}\\ y&\mapsto\big{(}\xi_{\alpha}(y),\pi(\xi_{\alpha}(y)^{-1}y)\big{)}\end{split}$
@@
(7.4.25) ${\mathrm{Gr}^{0}_{-\alpha}\times\mathrm{Gr}_{\mu+\alpha}}$${(\overset{\circ}{\mathfrak{G}}_{\alpha})_{\mu}}$$\scriptstyle{\mathbf{m}}$$\scriptstyle{F_{\alpha}}$
@@
We proceed by directly showing that $\mathbf{m}\circ F_{\alpha}={\rm id}_{\overset{\circ}{\mathrm{Gr}}_{\mu}}$ and $F_{\alpha}\circ\mathbf{m}={\rm id}_{\mathrm{Gr}^{0}_{-\alpha}\times\mathrm{Gr}_{\mu+\alpha}}$.
First, we check $\mathbf{m}\circ F_{\alpha}={\rm id}_{\overset{\circ}{\mathrm{Gr}}_{\mu}}$. Let $y\in\overset{\circ}{\mathrm{Gr}}_{\mu}$, then
@@
But, from Proposition 7.1$(iii)$ we see that $\xi_{\alpha}(y)n\xi_{\alpha}(y)^{-1}=n^{\prime}\in N^{\alpha}[t]$. Thus,
@@
Note that $y_{1}\in{\rm im}\,\tau_{\alpha}$; so by Proposition 7.1$(iii)$, we once again have that $y_{1}^{-1}ny_{1}=n^{\prime}\in N^{\alpha}[t]$. Therefore,
@@
The character $\chi_{\alpha}$ is contained in $\overset{\circ}{\mathfrak{G}}_{\alpha}$. By Proposition 7.11, we can identify its preimage under the moment map with $\mathbb{G}_{\alpha}\times\mathrm{Gr}_{\mu+\alpha}$. Moreover, by Theorem 7.8 this identification is $\mathbb{G}_{\alpha}$-equivariant. Thus,
(7.4.33) $\mathrm{Gr}_{\mu}/\!\!/\!\!/_{\chi_{\alpha}}\mathfrak{G}_{\alpha}\cong(\mathbb{G}_{\alpha}\times\mathrm{Gr}_{\mu+\alpha})/\mathbb{G}_{\alpha}\cong\mathrm{Gr}_{\mu+\alpha}~{}.$
@@
Let $\mu$ be a coweight of ${{\textup{PGL}}}_{n}$ and $\alpha$ a positive coroot, then the tuple $(\overset{\circ}{\mathfrak{G}}_{\alpha},\mathbf{m})$ is an inverse Hamiltonian reduction for $(\mathrm{Gr}_{\mu},\mathbb{G}_{\alpha},\Phi_{\alpha},\chi_{\alpha})$.
@@
(7.4.34) $\mathrm{Gr}_{\mu}^{\lambda}/\!\!/\!\!/_{\chi_{\alpha}}\mathfrak{G}_{\alpha}\cong\mathrm{Gr}_{\mu+\alpha}^{\lambda}~{}.$
Moreover, this reduction admits an inverse Hamiltonian reduction given by $(\overset{\circ}{\mathfrak{G}}_{\alpha},\mathbf{m})$.
@@
8. Inverse Hamiltonian reduction in type $A$
@@
Throughout this section, we fix $\mu$ to be a partition of $N$ of length $n-1$, thought of as a dominant coweight of $\rm PGL_{n}$ as in Section 3, and let $\alpha$ be a positive coroot of $\rm PGL_{n}$ such that $\mu+\alpha$ is again a dominant coweight, in keeping with the hypotheses of the preceding Section 7. Similarly, we return to the notational conventions of Section 3, for example letting $G={{\textup{GL}}}_{N}$ throughout.
8.1. Inverse Hamiltonian reduction for finite W-algebras
@@
are strongly $G$-equivariant with respect to the right $G$-actions on $S_{G,\mu+\alpha}$ and $S_{G,\mu}$, respectively, where we let $T^{*}\mathbb{G}_{\alpha}^{\textup{loc}}=\mathbb{G}_{\alpha}\times\overset{\circ}{\mathfrak{G}}_{\alpha}\cong\mathrm{Gr}_{-\alpha}^{0}$.
@@
The tuple, $(\overset{\circ}{\mathfrak{G}}_{\alpha},\mathbf{m})$ is an inverse Hamiltonian reduction for $(S_{G,\mu},\mathbb{G}_{\alpha},\Phi_{\alpha},\chi_{\alpha})$.
@@
Next, let us upgrade Corollary 8.3 to the quantum setting.
@@
Up to isomorphism, there exists a unique strongly $G\times\mathbb{G}_{\alpha}$ equivariant, graded quantization of the equivariant Slodowy slice $S_{G,\mu}$.
@@
The space of such quantizations are parameterized by the second $\mathbb{G}_{\alpha}\times G$-equivariant cohomology of $S_{G,\mu}$, by Proposition 4.3. Further, $H_{\mathbb{G}_{\alpha}\times G}^{2}(S_{G,\mu})=0$, noting $G$ acts freely with contractible quotient $S_{\mu}$, and $IG_{\alpha}$ is unipotent. Thus, there is a unique strongly $G\times\mathbb{G}_{\alpha}$-equivariant graded quantization, up to isomorphism. ∎
@@
The sheaf of algebras ${\mathcal{A}}_{G,\mu}$ is strongly $G\times\mathbb{G}_{\alpha}$-equivariant.
@@
This follows immediately from Lemmas 8.5 and 4.5, as ${\mathcal{A}}_{G,\mu}$ is the unique graded quantization of $S_{G,\mu}$, up to isomorphism. ∎
Combining both results, we have a novel proof of reduction by stages for $\hbar$-adic finite W-algebras in type $A$. This result was initially conjectured (in general type) by Morgan in [Mor15] and proven (likewise in general type) by Genra and Julliard in [GJ24], via homological methods.
@@
Let $\mu$ be a partition of $N$, thought of as a dominant coweight of $\rm PGL_{n}$, and $\alpha$ a positive root in such that $\mu+\alpha$ is dominant. Then, we have isomorphisms of graded quantizations.
@@
By Proposition 5.3, since $S_{G,\mu}$ admits an IHR, the Hamiltonian reduction of ${\mathcal{W}}_{\hbar,G,\mu}$ is a quantization of $S_{G,\mu+\alpha}$. However, from Lemma 4.5, there is a unique quantization of $S_{G,\mu+\alpha}$, up to isomorphism, and thus ${\mathcal{W}}_{\hbar,G,\mu}/\!\!/\!\!/_{\chi_{\alpha}}\cong{\mathcal{W}}_{\hbar,G,\mu+\alpha}$. Taking invariants with respect to the right $G$-action, we have the desired result for the ordinary (non-equivariant) finite W-algebras. ∎

### v2

@@ 7.3. Equivariance of the multiplication map
[…]
@@
[…]
@@
By Lemma 7.8(ii)(ii), $\mathbf{m}(y_{1},y_{2})=ny_{1}y_{2}n_{-}$ for some $n\in N^{\alpha}[t]$ and $n_{-}\in N_{-}[t]$. Therefore,
@@
But, from Proposition 7.1(iii)(iii), $x_{-\alpha}(\underline{v})nx_{-\alpha}(\underline{v})^{-1}=n^{\prime}\in N^{\alpha}[t]$. Putting things together,
@@ 7.4. Inverse Hamiltonian reduction for the translation group
From Lemma 7.8(iii)\,(iii), and Proposition 7.6, we see that the composition $\mathrm{Gr}^{0}_{-\alpha}\times\mathrm{Gr}_{\mu+\alpha}\xrightarrow{\bf m}\mathrm{Gr}_{\mu}\xrightarrow{\Phi_{\alpha}}\mathfrak{G}^{*}_{\alpha}$ factors through the inclusion $\overset{\circ}{\mathfrak{G}}_{\alpha}\hookrightarrow\mathfrak{G}^{*}_{\alpha}$. So, the image of the multiplication map lies inside the preimage of $\overset{\circ}{\mathfrak{G}^{*}_{\alpha}}$, i.e.,
(7.4.1) ${\bf m}:\mathrm{Gr}_{-\alpha}^{0}\times\mathrm{Gr}_{\mu+\alpha}\hookrightarrow(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha})_{\mu}~.$
@@
This follows immediately from Lemma 7.8(iii)(iii) and Proposition 7.7, since $\xi_{\alpha}$ is fully determined by the values of $\Phi_{\alpha}$ and $\zeta_{\alpha}$. ∎
So we see that $\xi_{\alpha}$ gives a putative section of ${\bf m}:\mathrm{Gr}_{-\alpha}^{0}\times\mathrm{Gr}_{\mu+\alpha}\hookrightarrow\Phi_{\alpha}^{-1}(\overset{\circ}{\mathfrak{G}_{\alpha}^{*}})_{\mu}$ to the first factor $\mathrm{Gr}_{-\alpha}^{0}$. We want to extend this to a full section.
@@
Let $y\in(\overset{\circ}{\mathfrak{G}^{*}_{\alpha}})_{\mu}$, then
@@
(7.4.24) $\begin{split}F_{\alpha}:(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha})_{\mu}&\rightarrow\mathrm{Gr}^{0}_{-\alpha}\times\mathrm{Gr}_{\mu+\alpha}\\ y&\mapsto\big{(}\xi_{\alpha}(y),\pi(\xi_{\alpha}(y)^{-1}y)\big{)}\end{split}$
@@
(7.4.25) $\mathrm{Gr}^{0}_{-\alpha}\times\mathrm{Gr}_{\mu+\alpha}$$(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha})_{\mu}$$\scriptstyle{\mathbf{m}}$$\scriptstyle{F_{\alpha}}$
@@
We proceed by directly showing that $\mathbf{m}\circ F_{\alpha}={\rm id}_{(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha})_{\mu}}$ and $F_{\alpha}\circ\mathbf{m}={\rm id}_{\mathrm{Gr}^{0}_{-\alpha}\times\mathrm{Gr}_{\mu+\alpha}}$.
First, we check $\mathbf{m}\circ F_{\alpha}={\rm id}_{(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha})_{\mu}}$. Let $y\in\overset{\circ}{\mathrm{Gr}}_{\mu}$, then
@@
But, from Proposition 7.1(iii)(iii) we see that $\xi_{\alpha}(y)n\xi_{\alpha}(y)^{-1}=n^{\prime}\in N^{\alpha}[t]$. Thus,
@@
Note that $y_{1}\in{\rm im}\,\tau_{\alpha}$; so by Proposition 7.1(iii)(iii), we once again have that $y_{1}^{-1}ny_{1}=n^{\prime}\in N^{\alpha}[t]$. Therefore,
@@
The character $\chi_{\alpha}$ is contained in $\overset{\circ}{\mathfrak{G}^{*}}_{\alpha}$. By Proposition 7.12, we can identify its preimage under the moment map with $\mathbb{G}_{\alpha}\times\mathrm{Gr}_{\mu+\alpha}$. Moreover, by Theorem 7.9 this identification is $\mathbb{G}_{\alpha}$-equivariant. Thus,
(7.4.33) $\mathrm{Gr}_{\mu}/\!\!/\!\!/_{\chi_{\alpha}}\mathbb{G}_{\alpha}\cong(\mathbb{G}_{\alpha}\times\mathrm{Gr}_{\mu+\alpha})/\mathbb{G}_{\alpha}\cong\mathrm{Gr}_{\mu+\alpha}~.$
@@
Let $\mu$ be a coweight of ${{\textup{PGL}}}_{n}$ and $\alpha$ a positive coroot, then the tuple $(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha},\mathbf{m})$ is an inverse Hamiltonian reduction for $(\mathrm{Gr}_{\mu},\mathbb{G}_{\alpha},\Phi_{\alpha},\chi_{\alpha})$.
@@
(7.4.34) $\mathrm{Gr}_{\mu}^{\lambda}/\!\!/\!\!/_{\chi_{\alpha}}\mathbb{G}_{\alpha}\cong\mathrm{Gr}_{\mu+\alpha}^{\lambda}~.$
Moreover, this reduction admits an inverse Hamiltonian reduction given by $(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha},\mathbf{m})$.
@@
8. Inverse Hamiltonian reduction in type AA
@@
Throughout this section, we fix $\mu$ to be a partition of $N$ of length n1n-1, thought of as a dominant coweight of $\rm PGL_{n}$ as in Section 3, and let $\alpha$ be a positive coroot of $\rm PGL_{n}$ such that $\mu+\alpha$ is again a dominant coweight, in keeping with the hypotheses of the preceding Section 7. Similarly, we return to the notational conventions of Section 3, for example letting $G={{\textup{GL}}}_{N}$ throughout.
8.1. Inverse Hamiltonian reduction for equivariant Slodowy slices
@@
are strongly $G$-equivariant with respect to the right $G$-actions on $S_{G,\mu+\alpha}$ and $S_{G,\mu}$, respectively, where we let $T^{*}\mathbb{G}_{\alpha}^{\textup{loc}}=\mathbb{G}_{\alpha}\times\overset{\circ}{\mathfrak{G}^{*}}_{\alpha}\cong\mathrm{Gr}_{-\alpha}^{0}$.
@@
The tuple, $(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha},\mathbf{m})$ is an inverse Hamiltonian reduction for $(S_{G,\mu},\mathbb{G}_{\alpha},\Phi_{\alpha},\chi_{\alpha})$.
@@
8.2. Existence of strong actions
We note that from [Los07], the quantizations ${\mathcal{A}}_{\hbar,G,\mu}$ are strongly $G$-equivariant for the right $G$-action on $S_{G,\mu}$. However, we must show that the Hamiltonian action of $\mathbb{G}_{\alpha}$ quantizes to a strong action of $\mathbb{G}_{\alpha}$ on the W-algebra.
@@ 8.2. Existence of strong actions
The action of $\mathbb{G}_{\alpha}$ on $S_{G,\mu}$ is free.
@@
Note that the action of $\mathbb{G}_{\alpha}$ on $S_{G,\mu}$ is the residual action of some unipotent subgroup of $G$ acting, via the left action, on the unreduced $T^{*}G$. Since this left action is free, this residual action on $S_{G,\mu}$ is free too. ∎
@@
The moment map $\Phi_{\alpha}:S_{G,\mu}\rightarrow\mathfrak{G}^{*}_{\alpha}$ is smooth.
@@
Note that the moment map $\Phi_{\alpha}$ is the base change of the same-named moment map $\Phi_{\alpha}:{\rm Gr}_{\mu}\rightarrow\mathfrak{G}^{*}_{\alpha}$ from Section 7.1, which is smooth by Lemma 7.3. Thus, $\Phi_{\alpha}:S_{G,\mu}\rightarrow\mathfrak{G}^{*}_{\alpha}$ is smooth too. ∎
These lemmas ensure that we can lift the Hamiltonian action.
Proposition 8.7.
Every, graded quantization of $S_{G,\mu}$ can be lifted to a strongly $\mathbb{G}_{\alpha}$-equivariant quantization.
In particular, there exists a quantum comoment map $\Phi_{\hbar,\alpha}:Sym(\mathfrak{G}_{\alpha}){{[\![\hbar]\!]}}\rightarrow{\mathcal{W}}_{\hbar,G,\mu}$, which integrates to an action of $\mathbb{G}_{\alpha}$. Moreover, every strongly $\mathbb{G}_{\alpha}$-equivariant quantization is isomorphic to ${\mathcal{W}}_{\hbar,G,\mu}$ and the isomorphism can be chosen to be strongly $\mathbb{G}_{\alpha}$-equivariant.
Proof.
From Lemmas 8.5 and 8.6, we note that the hypotheses of 4.6 are met. Thus, we have the desired lifitng property. The uniqueness follows from the fact that the moduli space of quantizations of $S_{G,\mu}$ is rigid from Lemma 4.8. ∎
As an immediate corollary, we note that ${\mathcal{A}}_{\hbar,G,\mu}$ is a strongly $\mathbb{G}_{\alpha}$-equivariant quantization of ${\mathcal{O}}_{S_{G,\mu}}$. We combine both results on equivariance in the following.
@@
Any graded, strongly $\mathbb{G}_{\alpha}\times G$-equivariant quantization of $S_{G,\mu}$ is isomorphic to ${\mathcal{W}}_{\hbar,G,\mu}$ and the isomorphism can be chosen to be strongly $\mathbb{G}_{\alpha}\times G$-equivariant.
Proof.
Note that ${\mathcal{W}}_{\hbar,G,\mu}$ is strongly $\mathbb{G}_{\alpha}\times G$-equivariant by Proposition 8.7 and [Los07]. Uniqueness follows from 4.5 and the fact that $H^{2}_{\mathbb{G}_{\alpha}\times G}(S_{G,\mu})=0$. ∎
Note that the restriction ${\mathcal{W}}_{\hbar,G,\mu}\hookrightarrow{\bf m}^{*}({\mathcal{W}}_{\hbar,G,\mu})$ is strongly $\mathbb{G}_{\alpha}$-equivariant, since the strong action on ${\bf m}^{*}({\mathcal{W}}_{\hbar,G,\mu})$ is defined by transport of structure.
Combining these observations, we have a version of reduction by stages for $\hbar$-adic finite W-algebras in type $A$. This result was initially conjectured (in general type) by Morgan in [Mor15] and proven (likewise in general type) by Genra and Juillard in [GJ24], via homological methods. Though, we note that our hypotheses on the nilpotent orbits are different to the ones in loc. cit.
Theorem 8.9.
We have a strongly $G$-equivariant isomorphism of strongly $G$-equivariant quantizations of $S_{G,\mu+\alpha}$:
@@
From Proposition 8.7, we note that ${\mathcal{W}}_{\hbar,G,\mu}$ has a quantum comoment map and thus we can make sense of the reduction. By Proposition 5.3, since $S_{G,\mu}$ admits an IHR, the Hamiltonian reduction of ${\mathcal{W}}_{\hbar,G,\mu}$ is a quantization of $S_{G,\mu+\alpha}$. However, from Lemma 4.8, there is a unique quantization of $S_{G,\mu+\alpha}$, up to isomorphism, and thus ${\mathcal{W}}_{\hbar,G,\mu}/\!\!/\!\!/_{\chi_{\alpha}}\cong{\mathcal{W}}_{\hbar,G,\mu+\alpha}$. Taking invariants with respect to the right $G$-action, we have the desired result for the ordinary (non-equivariant) finite W-algebras. ∎
8.3. Inverse Hamiltonian reduction for finite W-algebras